OpenStudy (anonymous):

cos(3x)=-1 using appropriate inverse trigonometric expression write and equation that defines the value of 3x solve this equation to find all possible values for angle 3x please help do not know where to start

OpenStudy (zzr0ck3r):

\[let \ \theta=3x\\cos(\theta)=-1\]what are the solutions for this?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

please help t

OpenStudy (zzr0ck3r):

then you need to memorize the unit cirle

OpenStudy (zzr0ck3r):

circle*

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

you should know all the values of x that make this true cos(x) = -1

OpenStudy (zzr0ck3r):

its at \[\{...-3\pi,-\pi,\pi,3\pi,.......\}\] so \[3x=\pi+2\pi*k \ \text{where} \ k\in \mathbb{Z}\]

OpenStudy (zzr0ck3r):

so k = {...-4,-3,-2,-1,0,1,2,3,4....}

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

np. memorize the unit circle. you really cant get by in this class without it.

OpenStudy (anonymous):

okay i will

OpenStudy (zzr0ck3r):

http://www.youtube.com/watch?v=OBmhgiLi77k this might help

OpenStudy (anonymous):

thank you!!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

this is the first part, with the inverse function \[3x=\cos^{-1}(-1)+2\pi k\]